MODS Micro LYNX Motor Controllers

This page reports the current status of the MODS Micro LYNX motor controllers. in response to problems that emerge.

Record changes here when...
  1. a mechanism cable is swapped to a spare controller after its controller fails or becomes marginal for a rapid on-telescope fix.
  2. a failed controller is replaced with a new controller.
  3. a new motor control code is uploaded into a controller.
Make changes in two places:
  1. Update the Mechanism Assignment Table below
  2. Add an entry to the Micro LYNX Change Record below. Entries are in reverse chronological order (new entires "push down" older entries).
For all reports please remember to date and initial entries, and where applicable note the relevant Issue Trak report (e.g., IT#1234) so that work recorded here in response to instrument issues can be tracked.

In your entries, make sure to explicitly state the instrument (MODS1 or MODS2), the channel (Red or Blue), the associated mechanism (e.g., rfilter, bcolttfa), and the controller ID (e.g., MLC15).

Micro LYNX Controller Assignments

Instrument Channel Controller Addr1 Port2 Mechanism Notes
MODS1 Red MLC1 101 8001 hatch  
    MLC2 101 8002 dichroic  
    MLC3 101 8003 spare (rgrating) moved from MLC6 2017 Oct 26 to run rgrating
unit failed 2018 Oct 28, rgrating cable moved to new MLC6
replaced 2019 Nov 27 (IT7935), now designated as spare w/rgrating code installed
R Pogge found it unresponsive on 2021 Feb 10; on 2/14/2021 P Hartley and B Bingham replaced it and R Pogge uploaded the rgrating code to it (IT 8317)
    MLC4 101 8004 rcolttfb  
    MLC5 101 8005 rcolttfc  
    MLC6 101 8006 rgrating unit failed 2017 Oct 26, replaced 8/2018 by Dan Pappalardo
moved rgrating here from MLC3 2018 Oct 29
unit was failing and was replaced 2019 Nov 27 (IT7935), remains rgrating
unit was again failing 2021 Feb 10, replaced 2/14/2021 (IT 8317) by P Hartley and B Bingham
unit failed once on 2022 May 24, replaced 2022 May 26 by C Bolyard
    MLC7 101 8007 rgrtilt1  
    MLC8 101 8008 rgrtilt2  
    MLC9 102 8009 rfilter  
    MLC10 102 8010 rcamfoc  
    MLC11 102 8011 rcolttfa moved from MLC3 2013 Sept (IT4801)
    MLC12 102 8012 minsert  
    MLC13 102 8013 mselect ext'd I/O board
    MLC14 102 8014 spare ext'd I/O board
    MLC15 102 8015 rshutter ext'd I/O board
    MLC16 102 8016 spare ext'd I/O board
  Blue MLC1 111 8001 calib  
    MLC2 111 8002 agwy  
    MLC3 111 8003 agwx  
    MLC4 111 8004 agwfoc  
    MLC5 111 8005 agwfilt  
    MLC6 111 8006 bcolttfa  
    MLC7 111 8007 bcolttfb  
    MLC8 111 8008 bcolttfc  
    MLC9 112 8009 bgrating unit failed 2020 Aug 27, replaced Aug 29 by Geno Bechetti. On 2021 Oct 07, it showed signs of failure and was replaced with a spare on 2021 Oct 08 by B Waggoner and J Prothro. That spare was then replaced by Chris Bolyard with another on 2021 Oct 28. Unit failing on 12/24/2022 (IT8794). JP swapped cable from MLC9 to MLC14. 4 Aug 2023, DP swapped in new spare, bgrating plc will need to be uploaded still. RP uploaded bgrating code on 8/18/23)
    MLC10 112 8010 bgrtilt1  
    MLC11 112 8011 bgrtilt2  
    MLC12 112 8012 bfilter  
    MLC13 112 8013 bcamfoc ext'd I/O board
    MLC14 112 8014 bgrating ext'd I/O board, was spare, being used for bgrating now (12/24/2022, IT8794). 4 Aug 2023, DP moved cable back to ML 9
    MLC15 112 8015 bshutter ext'd I/O board
    MLC16 112 8016 spare ext'd I/O board
MODS2 Red MLC1 201 8001 hatch  
    MLC2 201 8002 dichroic  
    MLC3 201 8003 rcolttfa  
    MLC4 201 8004 rcolttfb  
    MLC5 201 8005 rcolttfc  
    MLC6 201 8006 spare replaced 2018 Aug by Dan Pappalardo
    MLC7 201 8007 rgrtilt1  
    MLC8 201 8008 rgrtilt2 orig failed, replaced, reloaded rgrtilt2.plc 2016 Jan, started consistently to report TIMEOUT errors from 2017-01-06 (IT8795 & notes in IT5599); 2023 Aug 4, DanP swapped in a new spare, rgrtilt2 plc will need to be uploaded. RP uploaded rgrtilt2 code on 8/18/23.
    MLC9 202 8009 rfilter  
    MLC10 202 8010 rcamfoc  
    MLC11 202 8011 rgrating moved rgrating here 2015 April (IT5539)
    MLC12 202 8012 minsert  
    MLC13 202 8013 suspect Ext'd I/O board - was mselect, getting I/O read faults, swapped to MLC14 9/23/2018 (IT7296)
    MLC14 202 8014 mselect Ext'd I/O board - was spare, now mselect (9/23/2018 - IT7296)
    MLC15 202 8015 rshutter Ext'd I/O board
    MLC16 202 8016 spare Ext'd I/O board
  Blue MLC1 211 8001 calib  
    MLC2 211 8002 agwy  
    MLC3 211 8003 agwx  
    MLC4 211 8004 agwfoc  
    MLC5 211 8005 agwfilt  
    MLC6 211 8006 bcolttfa  
    MLC7 211 8007 bcolttfb  
    MLC8 211 8008 bcolttfc  
    MLC9 212 8009 bgrating  
    MLC10 212 8010 bgrtilt1  
    MLC11 212 8011 bgrtilt2  
    MLC12 212 8012 bfilter  
    MLC13 212 8013 bcamfoc Ext'd I/O board
    MLC14 212 8014 spare Ext'd I/O board
    MLC15 212 8015 bshutter Ext'd I/O board
    MLC16 212 8016 spare Ext'd I/O board

[1] Addr: base socket address is 192.168.139.

[2] Port: port assignment on the Comtrol RTS Eth-to-Serial Port server.

Micro LYNX Change Record

2023 Aug 4 [opk]

4-Aug-2023. During SSD with both MODS on their carts, on MODS1, D Pappalardo swapped the Blue Grat Sel mechanism cable from the spare ML 14 spot back to the ML 9 connector on the panel, labeled Grat Sel, and he opened the Blue IEB cover and replaced the bad ML 9 with a new spare from 3U [IT 8794]. On MODS2, he opened the red IEB cover and replaced the bad red IEB ML 8 with a new spare from 3U [IT 8795]. Once power and network connections are restored, Rick will need to upload the bgrating code to MODS1 Blue IEB ML 9 and the rgrating2tilt code to MODS2 Red IEB ML 8. The MODS1 mechanisms.ini file will need to be updated to re-assign the MODS1 bgrating select to Blue IEB port 9 instead of port 14.

18-Aug-2023. R Pogge uploaded bgrating code to MODS1 Blue IEB ML 9 and rgrating2tilt code to MODS2 Red IEB ML 8.

2022 Dec 24 [opk]

24-Dec-2022. MODS1 Blue Grating errors started on 24-Dec-2022 UT (IT8794). J Prothro & J Ornelas swapped the cable from MLC9 to MLC14, to which R Pogge uploaded the bgrating code. R Pogge updated the mechanisms.ini file on mods1 to re-assign the bgrating to MLC14. This is a temporary resolution and eventually MLC 9 should be replaced and the cable swapped back to MLC9.

2022 May [opk]

24-May-2021 MODS1 came up with the Red Grating showing no position data and a red background. Cycling power to MLC 6 resolved the immediate problem, though the microLynx controller was proactively replaced with a spare on 26-May-2022. Rick Pogge uploaded the rgrating code. Notes added to IT 8317.

2021 Oct [opk]

08-Oct-2021 MODS1 Blue Grating showed signs of failure (see IT8171). Elliott talked Brian Wagonner and Jay Prothro through removing the failing unit (MLC 9) and swapping in a spare. The spare was labelled Mar '21.

28-Oct-2021 Two more MODS1 bgrating timeout errors were seen (updates in IT8171) and Chris Bolyard replaced MLC 9 with another spare from the cabinet. Rick Pogge uploaded the bgrating code to it.

2021 Feb [opk]

MODS1 Red MLC6 running rgrating failed - see IT8317. It started to show signs of failure on 10 Feb 2021, but continued to work until 13 Feb 2021 when it failed again. It was replaced on 13 Feb 2021 by P Hartley and B Bingham, and R Pogge loaded the rgrating code into it.

MODS1 Red MCL3 which had been designated a spare and had the rgrating code uploaded was found unresponsive by R Pogge on 10 Feb 2021. It was replaced on 13 Feb 2021 by P Hartley and B Bingham, and R Pogge loaded the rgrating code into it.

2020 August [opk]

MODS1 Blue MLC9 running bgrating failed (serial comm unresponsive) - see IT8171. It was replaced on 28 Aug 2020 by G Bechetti and R Pogge loaded the bgrating code into it. This was the last spare, and OSU is shipping more tomorrow (31 Aug 2020).

2019 November [rwp]

MODS1 Red MLC6 running rgrating failed (serial comm unresponsive) - see IT7935. Replaced MLC6 and MLC3 (see 2018 August below). Loaded the rgrating code into MLC3 as a functional placeholder

2018 October [rwp]

MODS2 Red MLC3 running rgrating failed (serial comm unresponsive) - see IT6877. Moved rgrating to Red MLC6 and uploaded the rgrating code. Will be replaced at the same time as we replace Red MLC13.

2018 September [rwp]

MODS2 Red MLC13 (suspect slitmask select aka mselect) was giving apparent I/O sensor faults. Swapped to spare extended-I/O unit MLC14 and recovered proper mselect function (see IT7296). Red MLC13 is marked for replacement at the next convenient servicing opportunity.

2018 August [rwp]

MODS1 Red MLC6 (failed grating select) was not functioning, replaced with a spare unit by Dan Pappalardo, now designated as a normal-I/O spare without mechanism assignment.

MODS2 Red MLC6 showing amber LED on power-up, replaced with a spare unit by Dan Pappalardo, now designated as a normal-I/O spare without mechanism assignment.

2017 October [rwp]

MODS1 Red MLC6 (rgrating) failed on startup after reinstallation of MODS1 on the telescope. Swapped cable from Red MLC6 to Red MLC3 (unit replaced by OSU on Aug 2017), uploaded the rgrating.plc code, and changed the mechanism.ini file assignments. RGRATING now responsive, marking MLC6 as disabled in the table.

2017 August [rwp]

Replaced MODS1 Red MLC3 unit that failed on 2013 Sept 15 (IT#4801). Will load rcolttfa.plc code when MODS1 powered back on during Oct 2017 restart.

2017 June 6 [rwp]

Uploaded rgrating.plc into MODS2 Red MLC6 in anticipation that we might need to swap the red grating turret (rgrating) to this controller if the problem reported in IT#6643 becomes persistent and points to a problem with MODS2 Red MLC11 currently used to operate the MODS2 red grating turret (see 2015 Apr 24 below and IT#5539).

2016 January 5 [rwp]

Replaced MODS2 Red MLC6 (failed 2015 April 24, see below) with a new controller. While we had the MODS2 Red IEB open to replace this controller we noticed a fault indication (steady red LED lit) on MODS2 Red MLC8 (rgrtilt2), and replaced it as well. Perhaps not a coincidence we were also replacing the bad cable pin on rgrtilt2 during this service run.

Note that we left the MODS2 Red Grating Turret (rgrating) connected to MLC11 where it has been since the change reported below on 2015 April 24 (see IT#5539).

2015 April 24 [rwp]

MODS2 Red MLC6 failed (rgrating = red grating turret), reported in IT#5539. Swapped mechanism cable to MODS2 Red MLC11, uploaded rgrating.plc, and updated the IE mechanisms.ini file. For now we will leave the failed MODS2 Red MLC6 in place for future replacement.

2013 Sept 15 [rwp]

Problem emerged with the MODS1 red collimator TTF Actuator A (rcolttfa), reported in IT#4801, traced to failure of the MODS1 Red MLC3 unit. Solution was to perform a cable swap to MODS1 Red MLC11 and upload rcolttfa.plc into Red MLC11. For now we will leave the failed MODS1 Red MLC3 in place for future replacement.

-- %USERSIG{RichardPogge - 2017-07-26}%
Topic revision: r26 - 18 Aug 2023, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback